# A Practitioner’s Guide to Multi-turn Agentic Reinforcement Learning

### v1

@@ Abstract
We study what actually works and what doesn’t for training large language models as agents via multi-turn reinforcement learning. Despite rapid progress, existing frameworks and definitions are fragmented, and there is no systematic formulation or analysis of which design choices matter across tasks. We address this gap by first breaking down the design space into three inter-related pillars—environment, reward, and policy—and empirically derive a recipe for training LLM agents in situated textual domains. In particular, we test TextWorld and ALFWorld, popular domains for testing situated embodied reasoning, as well as SWE-Gym for more software engineering style tasks. (i) For the environment, we analyze the impacts of task complexity in terms of sizes of the state and action spaces as well as optimal solution length, finding that even simple environments within a domain can provide signal on how well an agent can generalize to more complex tasks. (ii) For the reward, we ablate relative reward sparsity, observing that while dense turn-level rewards accelerate training, performance and stability is highly dependent on the choice of RL algorithm. (iii) And for the agent’s policy, we explore the interplay between reward sparsity and biased (PPO, GRPO) and unbiased (RLOO) policy gradient methods in addition to showing how to find the optimal Supervised Fine-tuning (SFT) to RL training ratio given a fixed budget. We distill these findings into a training recipe that guides co-design across the three pillars, facilitating research and practical efforts in multi-turn agentic RL. Code: https://github.com/pearls-lab/meow-tea-taro
@@ 1 Introduction
Training LLMs as autonomous agents to navigate open-ended environments presents unique challenges: planning across extended horizons, making multi-turn sequential decisions, and optimizing for multi-turn rewards. The transition from static single-turn problem-solving to dynamic multistep reasoning is essential for agentic benchmarks such as interactive text and embodied simulations (TextWorld (Côté et al., 2018), ALFWorld (Shridhar et al., 2021), etc.), real-world software programming (OSWorld (Xie et al., 2024), SWE-gym (Pan et al., 2025), etc.), and abstract reasoning in novel situations (ARC-AGI (Chollet et al., 2025)). However, existing multi-turn RL implementations vary widely: some refer to tool-augmented single queries as multi-turn (Zeng et al., 2025), while many rely on model-based assumptions (Wang et al., 2025). This fragmentation has led to incomparable results across papers and confusion about what constitutes true multi-turn learning versus pseudo-multi-turn adaptations of single-turn methods.
@@
These insights yield a concrete multi-turn RL recipe that guides co-design across all three pillars. We demonstrate that multi-turn RL with LLMs is not simply an extension of single-turn optimization but requires fundamental rethinking across environment, policy, and reward. To facilitate future research, we will release our multi-turn agentic RL framework built on veRL (Sheng et al., 2025), including the training scripts and model weights of the experiments in this paper. This work provides both theoretical insights and practical guidelines for developing agentic AI systems capable of operating effectively in real-world interactive environments.
@@ 2 Related Work
While single-turn RL methods for LLMs including PPO (Schulman et al., 2017), RLOO (Ahmadian et al., 2024), GRPO (Shao et al., 2024), and DAPO (Yu et al., 2025) have been extensively optimized for immediate response quality, adapting them to multi-turn agentic scenarios remains non-trivial. These methods assume rewards directly follow individual actions, but multi-turn environments only reveal outcomes after extended interaction sequences, breaking the action-reward coupling that single-turn methods rely upon. Existing efforts on multi-turn RL have made limited progress on these challenges. Some approaches construct multi-turn scenarios by interleaving tool-use or reasoning steps for single-turn QA pairs (Zeng et al., 2025; Dong et al., 2025). Others who work on true interactive environments either rely on sparse terminal rewards without turn-level learning signals (Wang et al., 2025), or assign turn-level advantages uniformly across sequence tokens without fine-grained credit assignment (Zhou et al., 2025). More importantly, there lacks a comprehensive understanding of how the three fundamental pillars of RL – environment, policy, and reward – jointly determine performance in multi-turn interactive environments. This paper provides a systematic analysis on how the fundamental pillars of RL impact multi-turn RL training, respectively, and concludes insights on how to practically train multi-turn RL in different interactive agentic environments. Throughout the paper, we dedicate related works in individual sections.
@@ 3 Multi-turn Agentic Reinforcement Learning
We formulate multi-turn agentic tasks as a Partially Observable Markov Decision Process (POMDP) problem, defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\Omega,\mathcal{O},\gamma)$. Taking the Textworld task (Côté et al., 2018) as an example, an agent takes the action $a_{t}$ (go south) sampled from the action space $\mathcal{A}$ and receives a text observation $o_{t}$ (You are in front of a garden) from the observation space $\Omega$. $o_{t}$ is a partial description of the true state $s_{t}$ in the hidden state space $\mathcal{S}$ which contains the complete state world model. We assume that the state transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is deterministic. Upon taking an action, the agent also receives a scalar reward $r_{t}=\mathcal{R}(s_{t},a_{t})$. The agent’s objective is to learn a policy that maximizes the expected discounted sum of rewards $\mathbb{E}[\sum_{t}\gamma^{t}\cdot r_{t}]$.
@@ 4 Background and Experimental Setup
We evaluate on three situated textual benchmarks: TextWorld (Côté et al., 2018) and ALFWorld (Shridhar et al., 2021) for language grounding in situated embodied reasoning, and SWE-Gym (Pan et al., 2025) for real-world software engineering. We extend veRL (Sheng et al., 2025)’s efficient RL training infrastructure, integrating these benchmark environments through standard step and reset interfaces. Critically, unlike traditional RL settings that provide both observations and admissible action lists (reducing the problem to action selection), our agents must generate executable natural language commands from environment observations alone, without action hints—testing their ability to discover valid actions through exploration. The details of tasks and environments are provided in Appendix B.
@@
We use Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct [^2] https://huggingface.co/Qwen/Qwen2.5, and Qwen3-8B [^3] https://huggingface.co/Qwen/Qwen3-8B as base models (abbreviated as Qwen-1.5B, Qwen-7B, and Qwen-8B), training with PPO (Schulman et al., 2017), GRPO (Shao et al., 2024), and RLOO (Ahmadian et al., 2024) algorithms. Maximum iteration steps and token limits scale with task complexity. During rollout generation, we use temperature 0.7 to balance exploration-exploitation. We evaluate agents on held-out test sets, reporting task success rate as our primary metric—the percentage of episodes where agents achieve objectives within exploration budgets. For SWE-Gym, we report test suite passing ratios. Full training details and hyperparameter tuning analysis appear in Appendices C and A.
@@ 5 Environment
We design two experimental conditions to vary environment complexity and task diversity. For environment complexity, starting from the base configuration w2-o3-q4, we create controlled variations: w8-o3-q4 (increased spatial complexity), w2-o12-q4 (increased object complexity), w8-o12-q4 (increased both dimensions), and w4-o6-q8 (proportional scaling across all dimensions). For task diversity, we construct mixtures of increasing heterogeneity for training. In ALFWorld, we test single type (pick & place only), 4 types (heat, cool, cook, examine) and all 6 types (adding pick two & place). In SWE-Gym, we compare single type (getmoto only) against all 5 types (getmoto, pydantic, mypy, pandas, and iterative_dvc). All mixtures have the same data size to ensure fair comparison, with evaluation on both single-type and mixed-types held-out test sets. We train Qwen-1.5B, Qwen-7B, and Qwen-8B models using PPO and GRPO with consistent hyperparameters. Full task specifications and training details appear in Appendices B.1 and C.1.
@@ 5.1 How does multi-turn RL performance scale with environment complexity?
Table 1 reveals that base models struggle dramatically as environment complexity increases, with performance dropping from 17% to just 3% when both spatial and object dimensions are scaled. Critically, multi-turn RL improvements diminish with increasing complexity—while PPO achieves a 88% improvement on the base environment, this drops to 51% on the most complex setting. Notably, object complexity proves more challenging than spatial complexity, suggesting that manipulating and tracking multiple objects across turns presents fundamentally harder challenges than spatial exploration in situated textual domains.
When proportionally scaling all environment parameters, Table 2 shows that doubling all dimensions creates dramatically harder problems beyond linear scaling. The base Qwen-1.5B model’s performance collapses from 15% to 1%, indicating exponential search space expansion. While multi-turn PPO on Qwen-1.5B achieves substantial gains (58%), the final 59% success rate falls well short of the 80% achieved on w2-o3-q4. Performance also scales with model size—the 7B model reaches 72% success on w4-o6-q8, demonstrating that larger models better handle the increased state spaces of complex environments. Notably, even the 1.5B model shows strong learning potential with 65% and 58% gains on w2-o3-q4 and w4-o6-q8 respectively.
@@
Table 3: Multi-turn PPO performance on TextWorld w2-o3-q4 tasks with different exploration sizes.
#Exploration steps Qwen-1.5B Qwen-1.5B (PPO)
66 ($1.5\times$ optimal) $0.05$ $0.55_{\uparrow 0.5}$
88 ($2\times$ optimal) $0.09$ $0.73_{\uparrow 0.64}$
$12$ ($3\times$ optimal) $0.15$ $0.8_{\uparrow 0.65}$
$16$ ($4\times$ optimal) $0.17$ $0.88_{\uparrow 0.71}$
Lastly, we investigate how the exploration budget (the maximum steps agents can take during rollout) affects learning. For w2-o3-q4 tasks with 4-step optimal solutions, we vary maximum steps from 6 to 16. Table 3 shows that performance saturates beyond 8 exploration steps. Constraining agents to 6 steps ($1.5\times$ optimal) limits PPO to 55% success, while 8 steps ($2\times$ optimal) yields 73% success. Further increases to 12 and 16 steps produces only marginal gains. These results indicate that while insufficient exploration severely limits learning, budgets beyond 2× optimal provide negligible benefits for TextWorld tasks.
@@ 5.3 How does multi-turn RL generalize to different types of tasks within domain?
Table 5 and 6 show that multi-turn RL successfully applies to challenging ALFworld and SWE-Gym environments. Remarkably, agents trained on single task types achieve decent generalization to unseen task types, with 12% (ALFWorld) and 7% (SWE-Gym) improvements across all task types from single-type training alone. This indicates that multi-turn RL learns transferable skills even from limited task exposure. Surprisingly, multi-task training benefits seemingly unrelated tasks—agents trained on clean, heat, cool, and cook mixtures outperform single pick & place specialists by 19% on the single task itself and 21% on all-type evaluation. The results demonstrate that multi-turn RL develops generalizable skills that transfer across diverse objectives.
@@ 6 Policy
The choice of RL optimization algorithm and model initialization critically determines multi-turn RL performance. First, we examine the role of demonstration data. In practice, access to human demonstration data for supervised fine-tuning (SFT) may be limited. We ask: (1) What prior model policy enables effective multi-turn RL? This addresses whether expensive human demonstrations are necessary if agents can learn effectively from scratch. (2) Given a fixed data budget, what is the optimal ratio of SFT to RL data? This determines how to best allocate limited resources between demonstration collection and online learning. Second, we investigate whether RL optimization choices significantly impact multi-turn RL training. We compare heuristic policy gradient methods (PPO (Schulman et al., 2017), GRPO (Shao et al., 2024)) against unbiased methods (RLOO (Ahmadian et al., 2024)) to isolate the contributions of our multi-turn formulation from specific optimization heuristics—essential for making rigorous claims about algorithmic improvements, as demonstrated in recent work on the pitfalls of heuristic-dependent results (Oertell et al., 2025).
@@
TextWorld provides gold solutions for each procedurally generated game, which we use as demonstration data for SFT, representing human multi-turn trajectories. The SFT data follows a turn-based chat format, the default template used in most instruction-following scenarios. We generate SFT data with different random seeds than RL data to prevent leakage. We train Qwen-1.5B, Qwen-7B, and Qwen-8B models using PPO, GRPO, and RLOO with consistent hyperparameters across experiments. Full task specifications and training details appear in Appendices B.2 and C.2.
@@
For optimization algorithms with advantage estimation, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), we adopt token-level credit assignment. We compute token-level values and apply to TD error as $\delta_{t}^{i}=r_{t}^{i}+\gamma V(h_{t}^{i+1})-V(h_{t}^{i})$ where $h_{t}^{i}$ is the history up to and including token $a_{t}^{i}$. Then we estimate the advantage for each token using GAE: $\hat{A}_{t}^{i}=\sum_{l=0}^{L-i}(\gamma\lambda)^{l}\delta_{t}^{i+l}$, where $L$ is the horizon (number of tokens until episode ends). Even though only <eos> tokens receive rewards, all preceding tokens get non-zero advantages through value bootstrapping. Therefore, the Clipped Surrogate Objective for all tokens in the trajectory can be written as:

### v2

@@ Abstract
We study what actually works and what doesn’t for training large language models as agents via multi-turn reinforcement learning. Despite rapid progress, existing frameworks and definitions are fragmented, and there is no systematic formulation or analysis of which design choices matter across tasks. We address this gap by first breaking down the design space into three inter-related pillars—environment, reward, and policy—and empirically derive a recipe for training LLM agents in situated textual domains. In particular, we test TextWorld and ALFWorld, popular domains for testing situated embodied reasoning, as well as SWE-Gym for more software engineering style tasks. (i) For the environment, we analyze the impacts of task complexity in terms of sizes of the state and action spaces as well as optimal solution length, finding that even simple environments within a domain can provide signal on how well an agent can generalize to more complex tasks. (ii) For the reward, we ablate relative reward sparsity, observing that while dense turn-level rewards accelerate training, performance and stability is highly dependent on the choice of RL algorithm. (iii) And for the agent’s policy, we explore the interplay between reward sparsity and biased (PPO, GRPO) and unbiased (RLOO) policy gradient methods in addition to showing how to find the optimal Supervised Fine-tuning (SFT) to RL training ratio given a fixed budget.
@@ 1 Introduction
Training LLMs as autonomous agents to navigate open-ended environments presents unique challenges: planning across extended horizons, making multi-turn sequential decisions, and optimizing for multi-turn rewards. The transition from static single-turn problem-solving to dynamic multistep reasoning is essential for agentic benchmarks such as interactive text and embodied simulations (TextWorld (textworld), ALFWorld (alfworld), etc.), real-world software programming (OSWorld (osworld), SWE-gym (swegym), etc.), and abstract reasoning in novel situations (ARC-AGI (arcagi)). However, existing multi-turn RL implementations vary widely: some refer to tool-augmented single queries as multi-turn (two-turn-credit), while many rely on model-based assumptions (ragen). This fragmentation has led to incomparable results across papers and confusion about what constitutes true multi-turn learning versus pseudo-multi-turn adaptations of single-turn methods.
@@
These insights yield a concrete multi-turn RL recipe that guides co-design across all three pillars. We demonstrate that multi-turn RL with LLMs is not simply an extension of single-turn optimization but requires fundamental rethinking across environment, policy, and reward. To facilitate future research, we will release our multi-turn agentic RL framework built on veRL (verl), including the training scripts and model weights of the experiments in this paper. This work provides both theoretical insights and practical guidelines for developing agentic AI systems capable of operating effectively in real-world interactive environments.
@@ 2 Related Work
While single-turn RL methods for LLMs including PPO (ppo), RLOO (rloo), GRPO (grpo), and DAPO (dapo) have been extensively optimized for immediate response quality, adapting them to multi-turn agentic scenarios remains non-trivial. These methods assume rewards directly follow individual actions, but multi-turn environments only reveal outcomes after extended interaction sequences, breaking the action-reward coupling that single-turn methods rely upon. Existing efforts on multi-turn RL have made limited progress on these challenges. Some approaches construct multi-turn scenarios by interleaving tool-use or reasoning steps for single-turn QA pairs (two-turn-credit; arpo). Others who work on true interactive environments either rely on sparse terminal rewards without turn-level learning signals (ragen), or assign turn-level advantages uniformly across sequence tokens without fine-grained credit assignment (sweet-rl). More importantly, there lacks a comprehensive understanding of how the three fundamental pillars of RL – environment, policy, and reward – jointly determine performance in multi-turn interactive environments. This paper provides a systematic analysis on how the fundamental pillars of RL impact multi-turn RL training, respectively, and concludes insights on how to practically train multi-turn RL in different interactive agentic environments. Throughout the paper, we dedicate related works in individual sections.
@@ 3 Multi-turn Agentic Reinforcement Learning
We formulate multi-turn agentic tasks as a Partially Observable Markov Decision Process (POMDP) problem, defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\Omega,\mathcal{O},\gamma)$. Taking the Textworld task (textworld) as an example, an agent takes the action $a_{t}$ (go south) sampled from the action space $\mathcal{A}$ and receives a text observation $o_{t}$ (You are in front of a garden) from the observation space $\Omega$. $o_{t}$ is a partial description of the true state $s_{t}$ in the hidden state space $\mathcal{S}$ which contains the complete state world model. We assume that the state transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is deterministic. Upon taking an action, the agent also receives a scalar reward $r_{t}=\mathcal{R}(s_{t},a_{t})$. The agent’s objective is to learn a policy that maximizes the expected discounted sum of rewards $\mathbb{E}[\sum_{t}\gamma^{t}\cdot r_{t}]$.
@@ 4 Background and Experimental Setup
We evaluate on three situated textual benchmarks: TextWorld (textworld) and ALFWorld (alfworld) for language grounding in situated embodied reasoning, and SWE-Gym (swegym) for real-world software engineering. We extend veRL (verl)’s efficient RL training infrastructure, integrating these benchmark environments through standard step and reset interfaces. Critically, unlike traditional RL settings that provide both observations and admissible action lists (reducing the problem to action selection), our agents must generate executable natural language commands from environment observations alone, without action hints—testing their ability to discover valid actions through exploration. The details of tasks and environments are provided in Appendix B.
@@
We use Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct, and Qwen3-8B as base models (abbreviated as Qwen-1.5B, Qwen-7B, and Qwen-8B), training with PPO (ppo), GRPO (grpo), and RLOO (rloo) algorithms. Maximum iteration steps and token limits scale with task complexity. During rollout generation, we use temperature 0.7 to balance exploration-exploitation. We evaluate agents on held-out test sets, reporting task success rate as our primary metric—the percentage of episodes where agents achieve objectives within exploration budgets. For SWE-Gym, we report test suite passing ratios. Full training details and hyperparameter tuning analysis appear in Appendices C and A.
@@ 5 Environment
We design two experimental conditions to vary environment complexity and task diversity. For environment complexity, starting from the base configuration w2-o3-q4, we create controlled variations: w8-o3-q4 (increased spatial complexity), w2-o12-q4 (increased object complexity), w8-o12-q4 (increased both dimensions), and w4-o6-q8 (proportional scaling across all dimensions). In SWE-Gym, we focus on the getmoto task subset to isolate complexity factors, categorizing tasks into Easy, Medium, and Hard based on two distinct metrics: (1) Solution Complexity (lines of code in the gold patch) and (2) Verification Complexity (size of the test suite). For task diversity, we construct mixtures of increasing heterogeneity for training. In ALFWorld, we test single type (pick & place only), 4 types (heat, cool, cook, examine) and all 6 types (adding pick two & place). In SWE-Gym, we compare single type (getmoto only) against all 5 types (getmoto, pydantic, mypy, pandas, and iterative_dvc). All mixtures have the same data size to ensure fair comparison, with evaluation on both single-type and mixed-types held-out test sets. We train Qwen-1.5B, Qwen-7B, and Qwen-8B models using PPO and GRPO with consistent hyperparameters. Full task specifications and training details appear in Appendices B.1 and C.1.
@@ 5.1 How does multi-turn RL performance scale with environment complexity?
Table 1 reveals that base models struggle dramatically as environment complexity increases, with performance dropping from 17% to just 3% when both spatial and object dimensions are scaled. Critically, multi-turn RL improvements diminish with increasing complexity—while PPO achieves a 88% improvement on the base environment, this drops to 51% on the most complex setting. Notably, object complexity proves more challenging than spatial complexity, suggesting that manipulating and tracking multiple objects across turns presents fundamentally harder challenges than spatial exploration in situated textual domains. We observe parallel trends in the coding domain (Table 2). Base model performance degrades as tasks require longer code solutions or pass more rigorous test suites. However, Multi-turn GRPO consistently recovers performance across all difficulty tiers, validating that our metrics capture genuine complexity.
Performance also scales with model size. As shown in Table 3, the 7B model reaches 72% success on w4-o6-q8, demonstrating that larger models better handle the increased state spaces of complex environments. Notably, even the 1.5B model shows strong learning potential with 65% and 58% gains on w2-o3-q4 and w4-o6-q8 respectively.
@@
Table 2: Multi-turn GRPO performance on SWE-Gym tasks across varying complexity dimensions. We compare the Qwen3-8B base model against the policy trained with Multi-turn GRPO, categorizing tasks by solution size (lines of patch) and verification effort (test suite size).
Task Complexity Metrics Qwen3-8B (Base) Qwen3-8B (GRPO)
Metric A: Solution Complexity (Lines of Code)
Easy ($\leq$ 8 lines) $0.048$ $0.072_{\uparrow 0.024}$
Medium (9–31 lines) $0.048$ $0.109_{\uparrow 0.061}$
Hard (>> 31 lines) $0.018$ $0.096_{\uparrow 0.078}$
Metric B: Verification Complexity (Test Suite Size)
Easy ($\leq$ 17 tests) $0.040$ $0.133_{\uparrow 0.093}$
Medium (18–53 tests) $0.028$ $0.075_{\uparrow 0.047}$
Hard (>> 53 tests) $0.017$ $0.105_{\uparrow 0.088}$
@@
Table 4: Impact of exploration horizon on multi-turn performance. We compare the effect of increasing allowed exploration steps on TextWorld w2-o3-q4 tasks(Left) and allowed tool calls/response length on SWE-Gym tasks (Right).
TextWorld (PPO)
Model: Qwen-2.5-1.5B
Exploration Steps Base PPO
66 $0.05$ $0.55_{\uparrow 0.50}$
88 $0.09$ $0.73_{\uparrow 0.64}$
$12$ $0.15$ $0.80_{\uparrow 0.65}$
$16$ $0.17$ $0.88_{\uparrow 0.71}$
SWE-Gym (GRPO)
Model: Qwen3-8B
Tool Calls (Context) Base GRPO
$10$ (3k tokens) $0.03$ $0.03_{\uparrow 0.00}$
$25$ (4k tokens) $0.04$ $0.18_{\uparrow 0.14}$
$40$ (6k tokens) $0.04$ $0.22_{\uparrow 0.18}$
Lastly, we investigate how the exploration budget (the maximum steps agents can take during rollout) affects learning. For w2-o3-q4 tasks with 4-step optimal solutions, we vary maximum steps from 6 to 16. Table 4 shows that performance saturates beyond 8 exploration steps. Constraining agents to 6 steps ($1.5\times$ optimal) limits PPO to 55% success, while 8 steps ($2\times$ optimal) yields 73% success. Further increases to 12 and 16 steps produces only marginal gains. These results indicate that while insufficient exploration severely limits learning, budgets beyond 2× optimal provide negligible benefits for TextWorld tasks. In contrast, complex coding tasks benefit significantly from extended horizons. In SWE-Gym, increasing the context window and allowed tool calls from 10 to 40 boosts success rates from 3% to 22%. Unlike simple text games, reasoning in software repositories does not saturate quickly; rather, sustained exploration allows agents to iteratively debug and refine solutions.
@@ 5.2 How does multi-turn RL generalize to environments with different complexities?
This generalization capability extends to non-embodied domains. As shown in Table 6, code agents trained on ”Easy” tasks (based on test suite complexity) transfer effectively to harder problems, achieving gains of 4.8% on Medium and 3.6% on Hard tasks. This mirrors our findings in TextWorld, confirming that multi-turn RL acquires robust, reusable behaviors—such as error correction and state tracking—that remain effective even as the environment’s complexity scales.
@@
Table 6: Cross-environment generalization with Multi-turn GRPO on SWE-Gym tasks. We evaluate transfer performance across task complexity (Metric A: lines of patch) and verification complexity (Metric B: test suite size). Diagonal elements (gray) indicate in-distribution performance.
Metric A: Solution Complexity
Tasks easy medium hard
easy $072_{\uparrow 0.024} 0.0$$48_{\uparrow 0.0} 0.0$$028_{\uparrow 0.01}$
medium $0.161_{\uparrow 0.113}$ $109_{\uparrow 0.061} 0.0$$055_{\uparrow 0.037}$
hard $082_{\uparrow 0.034} 0.0$$91_{\uparrow 0.043} 0.0$$096_{\uparrow 0.078}$
Metric B: Verification Complexity
Tasks easy medium hard
easy $133_{\uparrow 0.093} 0.0$$76_{\uparrow 0.048} 0.0$$053_{\uparrow 0.036}$
medium $082_{\uparrow 0.042} 0.0$$75_{\uparrow 0.047} 0.0$$042_{\uparrow 0.025}$
hard $10_{\uparrow 0.06} 0.0$$062_{\uparrow 0.034}$ $0.105_{\uparrow 0.088}$
@@ 5.3 How does multi-turn RL generalize to different types of tasks within domain?
Table 7 and 8 show that multi-turn RL successfully applies to challenging ALFworld and SWE-Gym environments. Remarkably, agents trained on single task types achieve decent generalization to unseen task types, with 12% (ALFWorld) and 7% (SWE-Gym) improvements across all task types from single-type training alone. This indicates that multi-turn RL learns transferable skills even from limited task exposure. Surprisingly, multi-task training benefits seemingly unrelated tasks—agents trained on clean, heat, cool, and cook mixtures outperform single pick & place specialists by 19% on the single task itself and 21% on all-type evaluation. The results demonstrate that multi-turn RL develops generalizable skills that transfer across diverse objectives.
@@ 6 Policy
The choice of RL optimization algorithm and model initialization critically determines multi-turn RL performance. First, we examine the role of demonstration data. In practice, access to human demonstration data for supervised fine-tuning (SFT) may be limited. We ask: (1) What prior model policy enables effective multi-turn RL? This addresses whether expensive human demonstrations are necessary if agents can learn effectively from scratch. (2) Given a fixed data budget, what is the optimal ratio of SFT to RL data? This determines how to best allocate limited resources between demonstration collection and online learning. Second, we investigate whether RL optimization choices significantly impact multi-turn RL training. We compare heuristic policy gradient methods (PPO (ppo), GRPO (grpo), Reinforce++ (reinforce_pp)) against unbiased methods (RLOO (rloo)) to isolate the contributions of our multi-turn formulation from specific optimization heuristics—essential for making rigorous claims about algorithmic improvements, as demonstrated in recent work on the pitfalls of heuristic-dependent results (heuristics-harmful).
@@
TextWorld provides gold solutions for each procedurally generated game, which we use as demonstration data for SFT, representing human multi-turn trajectories. The SFT data follows a turn-based chat format, the default template used in most instruction-following scenarios. We generate SFT data with different random seeds than RL data to prevent leakage. We train Qwen-1.5B, Qwen-7B, and Qwen-8B models using PPO, GRPO, Reinforce++, and RLOO with consistent hyperparameters across experiments. Full task specifications and training details appear in Appendices B.2 and C.2.
@@
For optimization algorithms with advantage estimation, such as Proximal Policy Optimization (PPO) (ppo), we adopt token-level credit assignment. We compute token-level values and apply to TD error as $\delta_{t}^{i}=r_{t}^{i}+\gamma V(h_{t}^{i+1})-V(h_{t}^{i})$ where $h_{t}^{i}$ is the history up to and including token $a_{t}^{i}$. Then we estimate the advantage for each token using GAE: $\hat{A}_{t}^{i}=\sum_{l=0}^{L-i}(\gamma\lambda)^{l}\delta_{t}^{i+l}$, where $L$ is the horizon (number of tokens until episode ends). Even though only <eos> tokens receive rewards, all preceding tokens get non-zero advantages through value bootstrapping. Therefore, the Clipped Surrogate Objective for all tokens in the trajectory can be written as: